# Linear-quadratic optimal control for abstract differential-algebraic equations

### v1

@@ 5 Outlook
One interesting application of the presented approach is using the particular cost $\mathcal{Q}=0$, $\mathcal{R}=0$ and $\mathcal{N}=I$ which leads to the required supply $J(x_{0},u)=2\int_{0}^{T}\operatorname{Re}\langle y(t),u(t)\rangle{\rm d}t$. The Popov operator is then given by $\mathscr{R}:=\mathbb{F}+\mathbb{F}^{*}$. The coercivity of this operator can be characterized using the transfer operator. Furthermore, one can investigate possible relaxation of the assumption on the regularity and the index of the underlying differential-algebraic equation. For finite dimensional spaces $X$, $U$ and $Y$ there is the problem whether there exists a state feedback $F\in L(X,U)$ or output feedback $F\in L(Y,U)$ such that the closed loop pair $(E,A+BF)$ is regular with index one. For ordinary systems it is known that such a feedback exists if and only if a finite cost condition holds and it is expected that these results can be generalized to abstract differential-algebraic equations.

### v2

@@ 6 Outlook
One interesting application of the presented approach is using the particular cost $\mathcal{Q}=0$, $\mathcal{R}=0$ and $\mathcal{N}=I$ which leads to the required supply $J(x_{0},u)=2\int_{0}^{T}\operatorname{Re}\langle y(t),u(t)\rangle{\rm d}t$ and provides a link to passivity. The Popov operator is then given by $\mathscr{R}:=\mathbb{F}+\mathbb{F}^{*}$. Furthermore, one can investigate possible relaxation of the assumption on the regularity and the index of the underlying differential-algebraic equation. For finite dimensional spaces $X$, $U$ and $Y$ there is the problem whether there exists a state feedback $F\in L(X,U)$ or output feedback $F\in L(Y,U)$ such that the closed loop pair $(E,A+BF)$ is regular with index one. For ordinary systems it is known that such a feedback exists if and only if a finite cost condition holds and it is expected that these results can be generalized to abstract differential-algebraic equations.